An automobile tire is inflated to 35 psig at 54 oF. After being driven, the temperature rise to 80 oF. Determine the final gage pressure assuming volume remains constant


1
Expert's answer
2021-10-05T15:43:57-0400

Assuming ideal gas low

Process is Isochoric

"pV=nRT \\

pV = RT \\

\frac{p}{T} = \frac{R}{V} = constant \\

\frac{p_1}{T_1} = \frac{p_2}{V_2} \\

p_2 = \frac{p_1 \times T_2}{T_1} \\

T_1 = 54 =285 \;K \\

T_2 = 80 =299 \;K \\

p_2= \frac{35 \times 299}{285} = 36.72 \;psig"